Solve the differential equation (D^3+D^2+D+1)y=Sinx

1
Expert's answer
2022-02-08T12:08:13-0500

Corresponding homogeneous equation

"(D^3+D^2+D+1)y=0"

Characteristic (auxiliary)equation


"r^3+r^2+r+1=0"

"r^2(r+1)+(r+1)=0"

"(r+1)(r^2+1)=0"

"r_1=-1, r_2=i, r_3=-i"

The general solution of the homogeneous differential equation is


"y_h=c_1e^x+c_2\sin x+c_3\cos x"

The particular solution of the non homogeneous equation is


"y_p=Ax\sin x+Bx\cos x"

"y_p'=A\sin x+Ax\cos x+B\cos x-Bx\sin x"

"y_p''=2A\cos x-Ax\sin x-2B\sin x-Bx\cos x"

"y_p'''=-3A\sin x-Ax\cos x-3B\cos x+Bx\sin x"

Substitute


"-3A\sin x-Ax\cos x-3B\cos x+Bx\sin x"

"+2A\cos x-Ax\sin x-2B\sin x-Bx\cos x"

"+A\sin x+Ax\cos x+B\cos x-Bx\sin x"

"+Ax\sin x+Bx\cos x=\sin x"

"x\sin x:B-A-B+A=0"

"x\cos x:-A-B+A+B=0"

"\sin x:-3A-2B+A=1"


"\cos x:-3B+2A+B=0"


"A=B=-\dfrac{1}{4}"

The general solution of the given non homogeneous differential equation is


"y=c_1e^x+c_2\sin x+c_3\cos x-\dfrac{x}{4}\sin x-\dfrac{x}{4}\cos x"